OpenStudy (anonymous):

Any Help? " Imagine that you have $50 to spend at a used bookstore. You can buy paperbucks for $1 and hardbacks for $4. What combination of books might you buy with your $50?

jimthompson5910 (jim_thompson5910):

Let x = # of paperbacks y = # of hardbacks If paperbacks are $1 a piece, then buying x of them will cost 1*x = x dollars If hardbacks are $4 a piece, then buying y of them will cost 4*y = 4y dollars In total, you will spend x + 4y dollars If you set this total to be $50, then x+4y = 50

jimthompson5910 (jim_thompson5910):

Now let's say x = 2, if this is the case, then x+4y = 50 2 + 4y = 50 4y = 50-2 4y = 48 y = 48/2 y = 12 which means that you bought x = 2 paperbacks and y = 12 hardbacks

jimthompson5910 (jim_thompson5910):

What other combinations can you find?

OpenStudy (anonymous):

That Was All It Siad

jimthompson5910 (jim_thompson5910):

I know, but can you find any other combinations (other than x = 2, y = 12)

OpenStudy (anonymous):

I Really Don't Understand This ,At All That's Why I Asked You.. Sorry..

jimthompson5910 (jim_thompson5910):

what happens when x = 6? what is the value of y?

OpenStudy (anonymous):

Idk ..

jimthompson5910 (jim_thompson5910):

x+4y = 50 6+4y = 50 4y = 50-6 4y = 44 y = ??

OpenStudy (anonymous):

11

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so another combo is 6 paperbacks and 11 hardbacks

OpenStudy (anonymous):

So, that would be the answer?

jimthompson5910 (jim_thompson5910):

that's part of it, you can decide to find more combos if you want

jimthompson5910 (jim_thompson5910):

or you can stop here

OpenStudy (anonymous):

Okay... Thank You May I Ask U Something Else

jimthompson5910 (jim_thompson5910):

sure, what do you need?

OpenStudy (anonymous):

Write an equation in point-slope, slope-intercept, and standard forms for the line that is both parallel and perpendicular to the given line and that passes through the given point: y=-3x+5 (-3,1)

OpenStudy (anonymous):